OpenStudy (kyle755):

Solve: -1/2x+2=-x+7 A. 10 B. 5/2 C. 2/5 D. 20

jhonyy9 (jhonyy9):

repost it please on the mathematics topics section - thank you there will get help more easy sure

OpenStudy (jsidelnikov):

The Answer is A. What you do is you write down your problem and have -1/2x+2=-x+7 Then you move the variables or the x's to one side and the numbers in this case 2 to the other. You get:-1/2x+x=7-2. Then you take and find a common denominator which is easy so you have -1/2x+2x/2=7-2/ Then you add and get 1/2x=5 then you multiply by the reciprical of 1/2 which is 2 and have 5 times 2 which is 10 so the answer is A